clark earned $403 in simple interest on an investment of $5700. some of the money earned an interest rate of 4%, andthe rest earned 9%. How much did he invest at each rate?
Answer by elima(1433)

Let x be amount invested
.09x - amount invested at 9%
.04(5700-x) - amount invested at 4%
.09x + .04(5700-x)=403
.09x + 228-.04x=403
.09x-.04x=403-228
.05x=175
x=3500
amount invested at 9% = .09(3500) = 315
amount invested at 4% = .04(5700-3500) = .04(2200)= 88
:)